# Time is On My Side: Dynamics of Talk-Time Sharing in Video-chat Conversations

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing Collaborative and social computing design and evaluation methods
@@ 4.1. Preliminary measure: conversation-level imbalance
Table 2. Phrases that most distinguish between balanced and imbalanced conversations, and representative excerpts including them. The comparisons are done separately for comments discussing positive and negative aspects. Additional examples are included in Table 7 in the Appendix.
@@
Table 2 contains distinguishing phrases for highly-balanced and highly-imbalanced conversations, along with comment excerpts that contain them (additional examples in Table 7 in the Appendix). For a more interpretable analysis, we use the Fightin’ Words method to compare positive comments and negative comments separately. We see that speakers from highly-balanced conversations are more likely to refer to the joint aspects of the conversation: the most distinguishing words include “both”, “we both”, “our”, “each other”, and “lot in common”, perhaps suggesting connections between balance and other forms of discursive convergence (Guydish and Tree, 2023). On the other hand, speakers from highly-imbalanced conversations are more likely to refer to individuals (“he”, “she”, “me”, “i’m”, and “partner”), or to the skewed distribution of talk (“talked too much”). We find this joint-vs-individualistic distinction in both the positive and negative comments.
Figure 3. Percentage of primary and secondary speakers reporting maximum enjoyment scores for conversations with different levels of imbalance. The number of conversations for each level is indicated as $N$. Levels with 30 or less conversations ($N<30$) are not shown.
@@
This gap is also reflected when comparing the post-conversation comments of primary and secondary speaker in highly-imbalanced conversations, again using the Fightin’ Words method (Table 3, extended in Table 8 in the Appendix). Here, we find intuitive distinctions in the primary and secondary speakers’ characterizations of how much their conversation partners talked. Additionally, we reveal a dichotomy on how talkativeness (or lack thereof) is perceived. Secondary speakers can perceive the talkativeness of their partners positively, appreciating that they “had [a] lot of” interesting “stories” to “share”. But they can also perceive it as detrimental to the conversation flow: “it was hard” to contribute when the other “talked [a] lot”. This dichotomy is mirrored in the perceptions of the primary speakers: they can either feel connected when their partner is a “good listener” who shows “interest” by “asking questions”, or they can regret “talk[ing] too much” and perceive the other person as being too “shy”.
@@
Table 3. Phrases that distinguish most between comments by primary speakers and secondary speakers for highly-imbalanced conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments. Additional examples are included in Table 8 in the Appendix.
@@ 4.2. Main framework: talk-time sharing dynamics
One basic element of our formalism is a conversation window: an arbitrarily small span of the conversation covering a specific length of time. We take a sliding window approach to capture the way the conversation progresses: each conversation is composed of multiple windows of $K$ seconds, sliding in increments of $L$ seconds (and thus overlapping for $K-L$ seconds), where $L<K$ [^4] For our analysis, we use $K=150s$ and $L=30s$, arriving at these parameters after initial exploration of the CANDOR data; minor variations of $K$ and $L$ lead to similar results. For each such conversation window, we measure the imbalance of talk-time, i.e., the fraction of talk-time taken up by the most talkative speaker within that window (as opposed to across the entire conversation, as in Section 4.1). This simple approach allows us to capture the dynamics of talk-time sharing: at what moments is talk-time dominated by one speaker versus the other, at what moments is talk-time equally distributed, and in what sequence do these different regimes occur?
@@
A structured space of talk-time sharing dynamics. Together, the regimes depicted in Figure 4 amount to a fairly balanced conversation (overall conversation-level imbalance $=0.51$). However, in attending to the local dynamics of talk, we note that a different composition of regimes could have also resulted in the same conversation-level imbalance. For instance, a different, similarly-balanced conversation could consist of speakers who engage in more back-and-forths (i.e., more gray windows), rather than switching between longer stretches where one person leads.
@@
To further interpret the distinction between back-and-forth and alternating dominance conversations, we examine distinguishing phrases in the post-conversation comments, shown in Table 4 and Table 9 in the Appendix. Comparing positive comments, we note a stark difference in experience: while speakers in back-and-forth conversations are often only able to “keep the conversation going” by finding some common “topics”, those engaging in alternating dominance were more likely to be interested in “listen[ing]” to “the other’s” “stories” and “life experiences”. These two distinct ways of connecting with a stranger also came with their specific challenges. Many of the negative comments for back-and-forth conversations pointed to a necessary effortfulness: participants “had to” do something to “keep” the conversation “flow” and avoid awkward “moments”. The negative comments for alternating dominance were more diverse, though many questioned how genuine the interlocutor’s interest actually was. In summary, the differences in the talk-time sharing dynamics of balanced conversations pointed to distinctions between familiar modes of interaction—in this setting, sharing extended stories or self-disclosures (e.g., (Jefferson, 1978; Mandelbaum, 2012; Sprecher et al., 2013)) versus (perhaps awkwardly) searching for conversation topics (Jurafsky et al., 2009).
Table 4. Most distinguishing phrases between comments for stereotypical alternating dominance and back-and-forth conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments. Additional examples are included in Table 9 in the Appendix.
@@ A.1. Speaker Perception Across Axes
Figure 10. Similar to Figure 7, with similar percentage of windows that are neutral (gray $<5\%$), we show difference in speaker perception between conversations with more windows dominated by the secondary speaker (red) and conversations with more windows dominated by the primary speaker (blue).
@@
Figure 11. Similar to Figure 7, with similar percentage of windows dominated by the secondary speaker (red $<5\%$), we show difference in speaker perception between conversations with more windows in neutral (gray) and conversations with more windows dominated by the primary speaker (blue).
@@ A.5.1. More User Comment Examples
In this section, we provide expanded versions of Table 2,  3, and  4 including the full speaker comments to give a more complete picture of how users describe their experiences. See Table 7,  8, and  9.
Table 7. Extended version of Table 2, we show phrases that most distinguish between balanced and imbalanced conversations, and representative excerpts including them. The comparisons are done separately for comments discussing positive and negative aspects.
@@
Table 8. Extended version of Table 3, we show phrases that most distinguish between comments by primary speakers and secondary speakers for highly-imbalanced conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments.
@@
Table 9. Extended version of Table 4, we show most distinguishing phrases between comments for stereotypical alternating dominance and back-and-forth conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments.

### v3

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW330publicationmonth: 11doi: 10.1145/3757511ccs: Human-centered computing Collaborative and social computing design and evaluation methods
@@ 4.1. Preliminary measure: conversation-level imbalance
Table 2. Phrases that most distinguish between balanced and imbalanced conversations, and representative excerpts including them. The comparisons are done separately for comments discussing positive and negative aspects. Additional examples are included in Table 7 and  8 in the Appendix.
@@
Table 2 contains distinguishing phrases for highly-balanced and highly-imbalanced conversations, along with comment excerpts that contain them (additional examples in Table 7 and  8 in the Appendix). For a more interpretable analysis, we use the Fightin’ Words method to compare positive comments and negative comments separately. We see that speakers from highly-balanced conversations are more likely to refer to the joint aspects of the conversation: the most distinguishing words include “both”, “we both”, “our”, “each other”, and “lot in common”, perhaps suggesting connections between balance and other forms of discursive convergence (Guydish and Tree, 2023). On the other hand, speakers from highly-imbalanced conversations are more likely to refer to individuals (“he”, “she”, “me”, “i’m”, and “partner”), or to the skewed distribution of talk (“talked too much”). We find this joint-vs-individualistic distinction in both the positive and negative comments.
Figure 3. Percentage of primary and secondary speakers reporting maximum enjoyment scores for conversations with different levels of imbalance. The number of conversations for each level is indicated as NN. Levels with 30 or less conversations ($N<30$) are not shown.
@@
This gap is also reflected when comparing the post-conversation comments of primary and secondary speaker in highly-imbalanced conversations, again using the Fightin’ Words method (Table 3, extended in Table 9 and  10 inåthe Appendix).
Here, we find intuitive distinctions in the primary and secondary speakers’ characterizations of how much their conversation partners talked. Additionally, we reveal a dichotomy on how talkativeness (or lack thereof) is perceived. Secondary speakers can perceive the talkativeness of their partners positively, appreciating that they “had [a] lot of” interesting “stories” to “share”. But they can also perceive it as detrimental to the conversation flow: “it was hard” to contribute when the other “talked [a] lot”. This dichotomy is mirrored in the perceptions of the primary speakers: they can either feel connected when their partner is a “good listener” who shows “interest” by “asking questions”, or they can regret “talk[ing] too much” and perceive the other person as being too “shy”.
@@
Table 3. Phrases that distinguish most between comments by primary speakers and secondary speakers for highly-imbalanced conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments. Additional examples are included in Table 9 and  10 in the Appendix.
@@ 4.2. Main framework: talk-time sharing dynamics
One basic element of our formalism is a conversation window: an arbitrarily small span of the conversation covering a specific length of time. We take a sliding window approach to capture the way the conversation progresses: each conversation is composed of multiple windows of $K$ seconds, sliding in increments of $L$ seconds (and thus overlapping for KLK-L seconds), where L<KL<K. [^4] For our analysis, we use $K=150s$ and $L=30s$, arriving at these parameters after initial exploration of the CANDOR data; minor variations of $K$ and $L$ lead to similar results. For each such conversation window, we measure the imbalance of talk-time, i.e., the fraction of talk-time taken up by the most talkative speaker within that window (as opposed to across the entire conversation, as in Section 4.1). This simple approach allows us to capture the dynamics of talk-time sharing: at what moments is talk-time dominated by one speaker versus the other, at what moments is talk-time equally distributed, and in what sequence do these different regimes occur?
@@
A structured space of talk-time sharing dynamics. Together, the regimes depicted in Figure 4 amount to a fairly balanced conversation (overall conversation-level imbalance =0.51=0.51). However, in attending to the local dynamics of talk, we note that a different composition of regimes could have also resulted in the same conversation-level imbalance. For instance, a different, similarly-balanced conversation could consist of speakers who engage in more back-and-forths (i.e., more gray windows), rather than switching between longer stretches where one person leads.
@@
To further interpret the distinction between back-and-forth and alternating dominance conversations, we examine distinguishing phrases in the post-conversation comments, shown in Table 4 and Table 11 and  12 in the Appendix. Comparing positive comments, we note a stark difference in experience: while speakers in back-and-forth conversations are often only able to “keep the conversation going” by finding some common “topics”, those engaging in alternating dominance were more likely to be interested in “listen[ing]” to “the other’s” “stories” and “life experiences”. These two distinct ways of connecting with a stranger also came with their specific challenges. Many of the negative comments for back-and-forth conversations pointed to a necessary effortfulness: participants “had to” do something to “keep” the conversation “flow” and avoid awkward “moments”. The negative comments for alternating dominance were more diverse, though many questioned how genuine the interlocutor’s interest actually was. In summary, the differences in the talk-time sharing dynamics of balanced conversations pointed to distinctions between familiar modes of interaction—in this setting, sharing extended stories or self-disclosures (e.g., (Jefferson, 1978; Mandelbaum, 2012; Sprecher et al., 2013)) versus (perhaps awkwardly) searching for conversation topics (Jurafsky et al., 2009).
Table 4. Most distinguishing phrases between comments for stereotypical alternating dominance and back-and-forth conversations, and representative excerpts including them. The comparisons are done separately for positive comments and for negative comments. Additional examples are included in Table 11 and  12 in the Appendix.
@@ A.1. Speaker Perception Across Axes
Figure 10. Similar to Figure 7, with similar percentage of windows that are neutral (gray $5\%$), we show difference in speaker perception between conversations with more windows dominated by the secondary speaker (red) and conversations with more windows dominated by the primary speaker (blue).
@@
Figure 11. Similar to Figure 7, with similar percentage of windows dominated by the secondary speaker (red $5\%$), we show difference in speaker perception between conversations with more windows in neutral (gray) and conversations with more windows dominated by the primary speaker (blue).
@@ A.5.1. More User Comment Examples
In this section, we provide expanded versions of Table 2,  3, and  4 including the full speaker comments to give a more complete picture of how users describe their experiences. See Table 7,  8,  9,  10,  11, and  12.
Table 7. Extended version of Table 2: phrases from positive comments that most distinguish between balanced and imbalanced conversations, with representative excerpts.
@@
Table 8. Extended version of Table 2: phrases from negative comments that most distinguish between balanced and imbalanced conversations, with representative excerpts.
top distinctive phrases example excerpts
@@
Table 9. Extended version of Table 3: phrases that most distinguish between positive comments by primary and secondary speakers for highly-imbalanced conversations.
@@
Table 10. Extended version of Table 3: phrases that most distinguish between negative comments by primary and secondary speakers for highly-imbalanced conversations.
top distinctive phrases example excerpts
@@
Table 11. Extended version of Table 4: phrases that most distinguish between positive comments for stereotypical alternating dominance and back-and-forth conversations.
@@
Table 12. Extended version of Table 4: phrases that most distinguish between negative comments for stereotypical alternating dominance and back-and-forth conversations.
top distinctive phrases example excerpts